OpenStudy (anonymous):

Armando read 20 pages of his 200-page book in 4 hours. At this rate, how long will it take him to read the entire book?

OpenStudy (texaschic101):

20/4 = 200/x --- 20 pgs to 4 hrs = 200 pgs to x hrs cross multiply 20x = 800 -- divide by 20 x = 800/20 x = 40.....so it is 200 pgs in 40 hrs

OpenStudy (anonymous):

Can you help w some more?

OpenStudy (texaschic101):

I can try